I NEED TO KNOW FOR LINE SEGMENT AB WITH MIDPOINT M, DETERMINE THE COORDINATES OF POINT B. 46. A(15,25); M=(6.5, 16.5) CAN YOU SHOW ME STEP BY STEP HOW TO DO THIS? I DON'T UNDERSTAND THEM. THANK YOU.
Answer by stanbon(75887)

I NEED TO KNOW FOR LINE SEGMENT AB WITH MIDPOINT M, DETERMINE THE COORDINATES OF POINT B. 46. A(15,25); M=(6.5, 16.5) CAN YOU SHOW ME STEP BY STEP HOW TO DO THIS? I DON'T UNDERSTAND THEM. THANK YOU.
---------
Let B have coordinates: x,y
----
Since M is the middle: (15+x)/2 = 6.5
15 + x = 13
x = -2
-----
(25+y)/2 = 16.5
----
25+y = 33
y = 8
========
Cheers,
Stan H.
==========================